# MI-HET-855

### v1

@@ Appendix A $2\to 3$ cross section
$-1$ $\leq\cos{\theta^{\prime\prime}}\leq 1,$
00 0 $\leq\phi,\phi^{\prime\prime}\leq 2\pi,$
@@ Appendix C Kinematics of the single photon from $2\to 3$ scattering
Figure 4: MiniBooNE (left two panels) and MicroBooNE (right two panels) fits with neutrinos coupled to a $Z^{\prime}$ and $\phi$ exchanged with the nucleus where $m_{Z^{\prime}}=93.6$ MeV and $m_{\phi}=87.6$ MeV. The top panels show the fits only with neutrino scattering, while the bottom panels show the fits only with dark matter scattering where $m_{\chi}=m_{Z^{\prime}}/3$ and $g_{D}=0.5$. The neutrino fits require $g=g_{\nu}g_{\phi Z^{\prime}\gamma}y_{q}=9\times 10^{-7}\leavevmode\nobreak\ \text{GeV}^{-1}$, and the dark matter fits require $g=g_{\nu}g_{\phi Z^{\prime}\gamma}y_{q}=2\times 10^{-6}\leavevmode\nobreak\ \text{GeV}^{-1}$.
@@
    $T(-p_{23}^{*})$: The first is to boost the photon to the COM frame. This is done by boosting it by the momentum vector $-p_{23}^{*}$, where $p_{23}=p_{2}+p_{3}$ is given by
@@ Appendix D $\chi^{2}$ analysis
$dN^{\chi,{\rm sim}}_{\gamma}(m_{Z^{\prime}},m_{\phi},g)$ $=g^{2}\frac{N_{A}\rho_{T}l_{T}}{A_{T}}\sum_{E_{i}}N^{\chi}_{i}(m_{\chi}=m_{Z^{\prime}}/3,m_{Z^{\prime}})d\sigma(g_{\chi}=2.5,g_{\phi Z^{\prime}\gamma}=1,y_{q}=1,m_{\chi}=m_{Z^{\prime}}/3,m_{Z^{\prime}},m_{\phi}),$
$dN^{\nu,{\rm sim}}_{\gamma}(m_{Z^{\prime}},m_{\phi},g)$ $=g^{2}\frac{N_{A}\rho_{T}l_{T}}{A_{T}}\sum_{E_{i}}N^{\nu}_{i}d\sigma(g_{\nu}=1,g_{\phi Z^{\prime}\gamma}=1,y_{q}=1,m_{Z^{\prime}},m_{\phi}),$
where $d\sigma$ implies the total cross section, and where $N_{A}$ is the number of target nuclei inside the detector fiducial volume, $\rho_{T}$ and $A_{T}$ are the density (in g/cm3) and atomic number of the detector material, and $l_{T}$ is the depth of the detector (in cm).
Since more data are available from MiniBooNE’s neutrino mode, we fit our model to the excess observed in that mode. The best-fit parameters are those that minimize the quantity $(\chi^{2}/\text{dof})_{\text{cosine}}+(\chi^{2}/\text{dof})_{\text{energy}}$ or the MiniBooNE neutrino-mode excess. Using these best-fit parameters, we then calculate the predicted single-photon events at MicroBooNE. The $\chi^{2}$ statistic is defined as follows:
@@ Appendix E Fits
$\chi^{2}/$dof
No $(m_{\phi},m_{Z^{\prime}},m_{\chi})$ [MeV] $g$ [GeV-1] $\nu$-MB $\mu$B
1. (45.41, 48.49, 16.17) $5.7\times 10^{-7}$ 2.20 1.28
2. (87.67, 93.63, 31.21) $2.0\times 10^{-6}$ 1.89 1.27
3. (119.08, 114.06, 38.02) $3.3\times 10^{-6}$ 2.21 1.34
4. (193.07, 138.95, 46.31) $6.0\times 10^{-6}$ 1.79 1.41
5. (45.41, 48.49, 20.64) $5.7\times 10^{-7}$ 2.23 1.26
6. (87.67, 93.63, 39.8) $2.0\times 10^{-6}$ 1.94 1.28
7. (119.08, 114.06, 51.85) $3.2\times 10^{-6}$ 1.88 1.35
8. (193.07, 138.95, 87.7) $6.0\times 10^{-6}$ 1.80 1.40
@@
Table 2: Summary of MiniBooNE and MicroBooNE fits for various scalar and vector boson masses $m_{\phi}$ and $m_{Z^{\prime}}$ for dark matter scattering only (with mass $m_{\chi}$) scattering. The required combination of couplings is represented by $g=g_{\nu}g_{\phi Z^{\prime}\gamma}y_{q}$.
@@
Figure 4 shows the fits and corresponding $\chi^{2}$/dof values when considering the individual contribution from neutrinos and dark matter scattering. We observe that the fits closely resemble the combined fits presented in Fig. 2. Additionally, we report the $\chi^{2}$/dof values for the dark matter only fits, along with the required effective coupling values, assuming $\alpha_{\chi}=0.5$, in Table 2. We further find that the fits are not particularly sensitive to the choice of dark matter mass. This indicates that the dominant factors influencing the fits and $\chi^{2}$/dof values are the masses of the scalar and vector bosons.
Appendix F Bounds from charged meson three body decays
Figure 5: Bounds on $g_{\nu}$ as a function of $m_{Z^{\prime}}$ from three-body decay constraints set by PIENU and NA62. The fit points from Table 1 are represented by the blue circles.
Figure 5 depicts the bounds on $g_{\nu}$ as a function of $m_{Z^{\prime}}$. These constraints arise from the upper limit on the three body branching ratio of charged pions and kaons that are measured by the PIENU PIENU:2021clt and NA62 NA62:2021bji experiments respectively.

### v2

@@ Appendix A $2\to 3$ cross section
[…]
$0$ $\leq\phi,\phi^{\prime\prime}\leq 2\pi,$
@@ Appendix C Kinematics of the single photon from $2\to 3$ scattering
    $T(-p_{23}^{*})$: The first is to boost the photon to the COM frame. This is done by boosting it by the momentum vector $p_{23}^{*}$, where $p_{23}=p_{2}+p_{3}$ is given by
@@ Appendix D $\chi^{2}$ analysis
[…]
$dN^{\nu,{\rm sim}}_{\gamma}(m_{Z^{\prime}},m_{\phi},g)$ $=(g_{\nu}g_{\phi Z^{\prime}\gamma}y_{\phi})^{2}\frac{N_{A}\rho_{T}l_{T}}{A_{T}}\sum_{E_{i}}N^{\nu}_{i}d\sigma(g_{\nu}=1,g_{\phi Z^{\prime}\gamma}=1,y_{\phi}=1,m_{Z^{\prime}},m_{\phi}),$
where $d\sigma$ implies the total cross section, and where $N_{A}$ is the number of target nuclei inside the detector fiducial volume, $\rho_{T}$ and $A_{T}$ are the density (in g/cm3) and atomic number of the detector material, and $l_{T}$ is the depth of the detector (in cm). Also, $N_{i}^{\nu}$ and $N_{i}^{\chi}$ are the neutrino and dark matter fluxes reaching the detector, respectively. Here, we assume that the dark matter flux is derived after maximizing the branching ratio of the charged meson decay.
The best-fit parameters are those that minimize the quantity $(\chi^{2}/\text{dof})_{\text{cosine}}+(\chi^{2}/\text{dof})_{\text{energy}}$ or the MiniBooNE neutrino-mode excess. Using these best-fit parameters, we then calculate the predicted single-photon events at MicroBooNE. The $\chi^{2}$ statistic is defined as follows:
@@ Appendix E Fits
$g_{\phi Z^{\prime}\gamma}y_{\phi}$ $\chi^{2}/$dof
[…]
1. (45.41, 48.49, 16.17) $\times 10^{-4} 2.2$0 1.28
2. (87.67, 93.63, 31.21) $5\times 10^{-4}$ 1.89 1.27
3. (119.08, 114.06, 38.02) $7\times 10^{-4}$ 2.21 1.34
4. (193.07, 138.95, 46.31) $1\times 10^{-3}$ 1.79 1.41
5. (45.41, 48.49, 20.64) $\times 10^{-4} 2.2$3 1.26
6. (87.67, 93.63, 39.8) $5\times 10^{-4}$ 1.94 1.28
7. (119.08, 114.06, 51.85) $7\times 10^{-4}$ 1.88 1.35
8. (193.07, 138.95, 87.7) $1\times 10^{-3}$ 1.80 1.40
@@
Table 2: Summary of $\chi^{2}$ values for the MiniBooNE fits and MicroBooNE overlays for various values of $m_{\phi}$ and $m_{Z^{\prime}}$ for dark matter scattering only (with mass $m_{\chi}$). The two table sections correspond to different ratios of $m_{\chi}/m_{Z^{\prime}}$.
Figure 4: Fits to the MiniBooNE neutrino-mode (left), and overlays with anti-neutrino mode (middle), and the inclusive single photon MicroBooNE (right) cosine (upper) and visible energy (lower) excesses for ${\rm neutrino}+{\rm dark}$ matter scattering where the mediators used are $m_{Z^{\prime}}=93.6$ MeV, and $m_{\phi}=87.6$ MeV, and the combination of couplings are $g=g_{\nu}g_{\phi Z^{\prime}\gamma}y_{\phi}=8.2\times 10^{-7}\leavevmode\nobreak\ \text{GeV}^{-1}$. The black data points+error bars denote the statistical errors, and the colored error bars correspond to the systematic uncertainties.
@@
Figure 4 shows the MiniBooNE neutrino mode fits and overlays for the MiniBooNE antineutrino mode and MicroBooNE neutrino mode with the combined neutrino+dark matter signals. We observe that the fits closely resemble the individual fits presented in the paper. Additionally, we report the $\chi^{2}$/dof values for the dark matter only fits, along with the required effective coupling values, assuming $\alpha_{\chi}=0.5$, in Table 2. We further find that the fits are not particularly sensitive to the choice of dark matter mass. This indicates that the dominant factors influencing the fits and the values of $\chi^{2}$ / dof are the masses of the scalar and vector bosons.
Appendix F Coherent contribution
Figure 5 depicts the energy-cosine distribution of the photons from the coherent part of the $2\to 3$ process, where the average cross section is $\sigma_{\text{coh}}^{\text{avg}}\simeq 1\times 10^{-40}\leavevmode\nobreak\ \text{cm}^{2}$. This includes contributions from both neutrinos and dark matter. We see that this differs from the template used in Ref. MicroBooNE:2025rsd , with about $\sim 10\%$ overlap between them. Including this fact, we see that the cross section satisfies the upper bound on the coherent cross section.
Figure 5: The energy cosine distribution of the coherent $2\to 3$ process from neutrinos and dark matter for $m_{\phi}=87\leavevmode\nobreak\ \text{MeV},\leavevmode\nobreak\ m_{Z^{\prime}}=93\leavevmode\nobreak\ \text{MeV}$, and $g=8.2\times 10^{-7}\leavevmode\nobreak\ \text{GeV}^{-1}$.
Figure 6: Bounds on $y_{\phi}$ as a function of $m_{\phi}$ for a visible scalar, assuming $g_{Z^{\prime}}=2\times 10^{-2}$, $x_{u}=1,x_{d}=-1.2,x_{c}=-0.3,$ and $x_{s}=-1.016$. from BaBar and CHARM monophoton searches, for different mass splittings $\Delta m$ (see text for more details). The green dashed line shows the predictions for Belle-II at 50 ab-1.
Appendix G Bounds on visible decaying scalars
Figure 6 depicts the bounds on the Yukawa coupling $y_{\phi}$ for a visibly decaying scalar. If $y_{\chi}=0$, i.e., the scalar does not couple to dark matter, the scalar is rendered visible and can produce a single photon via $\phi\to Z^{\prime}\gamma$/ $\phi\to\bar{\chi}\chi\gamma$. For a visibly decaying scalar, bounds from CHARM are also applied. The decay width of a scalar is larger for a larger mass gap, where we define $\Delta m=m_{\phi}-m_{Z^{\prime}}$ or $m_{\phi}-2m_{\chi}$. We find that the bounds shift towards lower Yukawas for larger mass gaps due to lower lifetimes for larger mass gaps. Additionally, we see that the bounds from E949/E787 E787:2002qfb ; E787:2004ovg ; E949:2007xyy ; BNL-E949:2009dza have an additional upper limit, as larger couplings result in decays inside the detector, which does not render the scalar as “invisible”.